OpenStudy (anonymous):

Find the cube roots of 8(cos 216° + i sin 216°). Please help!

OpenStudy (anonymous):

Hope I can help. we have formula for this : let z = 8(cos 216 + i sin 216). now we apply formula z^1/3 = [8(cos 216 + isin216)]^1/3 = 8^1/3 (cos(216/3) + i sin (216/3)= 2 (cos 72 + i sin 72).